OpenStudy (anonymous):

A drawer contains 4 red socks, 3 white socks, and 3 blue socks. Without looking, you select a sock at random, replace it, and select a second sock at random. What is the probability that the first sock is blue and the second sock is red?

OpenStudy (anonymous):

A: 3/5 B: 7/20 C: 3/25 D: 7/100

OpenStudy (anonymous):

C: 3/25 P (blue) = 3/10 P ( red) = 4/10 ( with replacement) => P (blue then red) = 3/10 * 4/10 = 12/100 = 3/25

OpenStudy (anonymous):

OMG! Thanks Ur A Life Saver! (:

OpenStudy (anonymous):

Should you happy with it! Why sad face ???